Solve each system of equation
3x+2y=12
x+y=3
This problem is really confusing me I tried changing them in y=mx+b but the answer still isn't right and then I tried multiplying the bottom equation by -3 and I get another wrong answer. Please help.

Answer by richwmiller(17219)

addition/elimination
3x+2y=12
x+y=3
3x+2y=12
-3x-3y=-9
add
-y=3
y=-3
x+y=3
x-3=3
x=6
substitution
3x+2y=12
x+y=3
y=3-x
3x+2(3-x)=12
3x+6-2x=12
x=6
x+y=3
6+y=3
y=-3